Reports 1-1 of 1 Clear search Modify search
tomotada.akutsu - 22:04 Tuesday 02 March 2021 (16126) Print this report
Replacing the oplev TX holder for MCi

Replaced the oplev TX holder for MCi. Replacement itslef was easily done.

  1. The TX holder was replaced with the newer one. See Fig. 1 and Fig. 2.
  2. I roghly align the TX so that the oplev beam can come back to the QPD (oplev RX).
  3. By the way, I should have been more careful. When I checked the oplev-related signals in the site, I found beat notes in the MCi oplev (see Fig. 3) and got surprised; the faster component seems 8 cycles in 10 sec, so 0.8Hz, but not found in the spectrum. It turns out this is due to a large 120 Hz alias, and so it is not seen in the properly anti-aliased channel (Fig. 5) and higher-speed channel (Fig. 6; green is of today while blue is of yesterday). Anyway, such a large 120 Hz peak was not observed one day ago (see Fig. 4 as a comparison to Fig. 3; also see Fig. 6). So this might be related to my work; maybe related to that I turned on room lights and left as was.
  4. Anyway, at the site, I had done optical alignment several times!! to "recover" the non-beat note signals. During this hopeless work, I learned that the accessibility to MCi was very hard due to some beam dumps in front of MCi. Maybe detaching them once would be helpful...though I do not want to do so. It is hard to align the TX together with looking at the inside of MCF via the K400 flange, which was detached the other day (16120).
  5. Sorry, but I left the place without fixing the oplev TX and RX. Please never kick them off...

Memo: Because the oplev reference was or will be lost, it might be nice to input the laser light routinely to check whether the flashing can be observed, until the oplev will be fixed.

Looking at Figs. 3 and 4, the MCe oplev lost the light.

Images attached to this report
Comments to this report:
tomotada.akutsu - 17:24 Wednesday 03 March 2021 (16133) Print this report

Today I just turned off the room light in the MCF clean booth at 10:07 JST (01:07 UTC), and the result supports my estimation as shown in the attached figure.

Images attached to this comment
Search Help